A country with a population of eight million adults has five million employed, 500,000 unemployed, and the rest of the adult population is out of the labor force. What’s the unemployment rate? What share of population is in the labor force? Sketch a pie chart that divides the adult population into these three groups.
Answer #1

population adults = 8M M = millions

employed = 5M   

unemployed = 0.5M

labor force = 5 + 0.5 M = 5.5 M

unemployment rate = 0.5/ 5.5 M = 0.090909

share 0f population in labor force = 5.5/ 8 = 0.6875